OpenStudy (anonymous):

1.∫(2u+10)/(3u+13) du 2.∫(u+5)/(5u-2)^2 du

OpenStudy (anonymous):

\[I = \int\limits{\frac{u+5}{(5u-2)^2}}du = \int\limits{\frac{A}{(5u-2)^2 } + \frac{B}{(5u-2)}}du \] A + B(5u-2) = u+5 let u = 2/5 A = 27/5

OpenStudy (anonymous):

let u = 0 27/5 -2B = 5 B = 1/5

OpenStudy (lalaly):

For the first one do long division first

OpenStudy (anonymous):

eigen : A + B(5u-2) = u+5 why not A(5u-2)^2 + B(5u-2) =u+5 lalaly : complete answer pls I'm too slow --"

OpenStudy (lalaly):

Do u know how to do long divisio?

OpenStudy (anonymous):

if \[\frac{u+5}{(5u-2)^2} = \frac{A}{(5u-2)^2} + \frac{B}{5u-2}\] multiply both sides by the LHS denominator to get the equation i had

OpenStudy (wasiqss):

lana tell him the easier way out split into 2u/(10u+13) + 10/(10u+13)

OpenStudy (lalaly):

Ok but still how would u integrate the first one @wasiqss

OpenStudy (lalaly):

2u/(10u+13)?

OpenStudy (wasiqss):

yes lana imma tell you

OpenStudy (lalaly):

Either way u should do long division @wasiqss

OpenStudy (wasiqss):

hmm partial fracction, but yehh it will lead us to same route

OpenStudy (wasiqss):

yayyyyyyy i found the easier method

OpenStudy (wasiqss):

multiply divide by 5

OpenStudy (wasiqss):

take 1/5 as common

OpenStudy (anonymous):

lalaly's solution (i think is fairly simple anyway)  wasiqss: 2u/(10u+13) will integrate by parts, not really worth it though